# Integrated modeling of boron powder injection for real-time plasma-facing component conditioning

### v1

@@ Abstract
An integrated modeling framework for investigating the application of solid boron (B) powder injection for real-time surface conditioning of plasma-facing components (PFCs) in tokamak environments is presented. Utilizing the DIII-D impurity powder dropper (IPD) setup, this study simulates B powder injection scenarios ranging from milligrams to tens of milligrams per second, corresponding to boron flux rates of $10^{20}-10^{21}$ B/s in standard L-mode conditions. The comprehensive modeling approach combines EMC3-EIRENE for simulating the plasma background and the Dust Injection Simulator (DIS) for the ablation and transport of the boron powder particles. The results show substantial transport of boron to the inboard lower divertor, predominantly influenced by the main ion plasma flow. The powder particle size (5-250 $\mu$m) was found to be insignificant for the scenario considered. The effects of erosion and redeposition were considered to reconcile the discrepancies with experimental observations, which saw substantial deposition on the outer plasma-facing components. For this purpose, the WallDYN3D code was updated to include boron sources within the plasma domain and integrated into the modeling framework. The mixed-material migration modeling shows evolving boron deposition patterns, suggesting the formation of mixed B-C layers or predominantly B coverage depending on the powder mass flow rate. While the modeling outcomes at lower B injection rates tend to align with DIII-D experimental observations, the prediction of near-pure boron layers at higher rates has yet to be experimentally verified in the carbon environment of the DIII-D tokamak. The extensive reach of boron layers found in the modeling suggests the need for modeling that encompasses the entire wall geometry for more accurate experimental correlations. This integrated approach sets a precedent for analyzing and applying real-time in-situ boron coating techniques in advanced tokamak scenarios, potentially extendable to ITER.
@@ 1 Introduction
The technique of real-time wall conditioning and coating (RTWC) became recently subject of numerous studies in tokamaks and stellarators [1, 2, 3, 4, 5, 6, 7, 8, 9]. Traditional wall conditioning (boronization) techniques involve the injection of hazardous borane gases and interrupting plasma operation, falling short in the context of long-pulse steady-state reactor operation. Moreover, they often fail to condition the divertor plasma-facing components (PFCs) as efficiently as the main chamber walls. For ITER, the strategic decision to replace the beryllium wall with a full-tungsten (W) one necessitates adjustments to the baseline scenario and additional risk-mitigation strategies to sustain high-plasma performance, especially during the critical Q=10 phase. In this context, investigating alternative methods like boron material injection could be beneficial. The evolution of RTWC research has been greatly facilitated by the development of the impurity powder dropper (IPD) technology, designed for releasing single or multi-species impurity powders into the plasma in real-time [10]. The IPD can dispense a wide range of impurities in a non-gaseous state, including but not limited to lithium, boron, boron nitride, and silicon. Comprehending the transport and deposition processes of locally released impurities in the scrape-off-layer (SOL) holds key importance in developing this technology for ITER and next-generation fusion devices.
In this paper, a new modeling framework for injecting solid boron, the primary material used for real-time wall conditioning in current tokamaks, stellarators, and heliotrons, will be presented. It introduces a comprehensive, integrated approach to model the transport, erosion, and deposition of boron powder, illuminating the extent of coverage achieved on the main PFCs. This will be crucial for evaluating the suitability of this technique for next-generation fusion devices like ITER and Fusion Power Plants (FPPs), which require novel solutions to maintain wall integrity and sustain high-performance plasmas during extended operations [11, 12, 13].
@@
The EMC3-EIRENE code [14, 15] effectively simulates full torus plasma, impurity transport, and plasma-material interactions. Previous studies using this code have identified non-axisymmetric effects from atomic impurity sources in axis-symmetric configurations, such as nitrogen gas seeding in Alcator C-Mod and boron injection in DIII-D [16, 17]. The Dust Injection Simulator (DIS) code [18] is used to model dust transport, similar to the DUSTT code, which has been applied to intrinsic dust in 2D plasma backgrounds [19, 20] and injected dust in the 3D heliotron LHD [21].
@@
To decipher the mechanisms that lead to B wall coatings through B powder injection, it is crucial to correlate deposition with transport. For this purpose, the WallDYN3D code was employed, which has been recently extended to 3D configurations to enable studies in both tokamaks with non-axisymmetric wall geometries, RMPs, or other symmetry breaking such as local gas puffs and inherently three-dimensional configurations like stellarators [22, 23, 24].
Figure 1: (a) Top-down view and cross-sectional view on boron source distribution resulting from ablation of 5-250 $\mu$m size powder particles along there trajectories calculated with DIS. (c) Boron flux at peak location (inner divertor) for 5-250 $\mu$m size boron powder particles originating from B powder injection (zero impurity recycling).
The forthcoming sections of this paper briefly introduce the three models used (Sec. 2 and 3), discuss the transport effects, and resulting B distribution without impurity recycling (Sec. 4), proceed with the analysis based on mixed-material modeling (Sec. 3 and 6), and finally, present a summary and conclusions (Sec. 7).
@@ 2 Coupled dust and plasma impurity transport modeling with DIS and EMC3-EIRENE
The Dust Injection Simulator (DIS) model was designed to study the dynamics of dust particles within three-dimensional, time-evolving plasmas [18]. It employs a comprehensive approach, integrating equations that account for the motion, charge balance, temperature changes, and mass evolution of the dust particles. These equations allow for the simulation of dust transport and ablation processes, including sublimation.
@@
where $R_{f}$ describes the force acting on the powder particles consisting of the drag force exerted by the plasma ($F_{drag}$), electrostatic force ($F_{E}$), gravitational force ($F_{grav}$), and centrifugal force ($F_{centr.}$). Presently, $F_{E}$ does not contribute to the dust transport since EMC3-EIRENE does not yet solve the electric field self-consistently. The deposition depths in the plasma of the injected materials are determined by material-specific properties such as vapor pressure, heat sublimation, emissivity, secondary electron emission, material density, and boiling and melting points. Note that the present version of the model assumes perfect spherical symmetry of particles, neglecting effects such as the rocket force [25].
@@
where $S_{z}$ denotes the impurity ionization source for charge state $Z$, and $D_{imp}$ the anomalous particle diffusivity. The collision time $\tau_{zi}$ between impurities and background main ions is calculated as $\tau_{zi}=\frac{1.47\cdot 10^{13}m_{z}(\nicefrac{{T_{i}^{3}}}{{m_{i}}})^{0.5}}{(1+\nicefrac{{m_{i}}}{{m_{z}}})n_{i}Z^{2}ln\Lambda}$ ($ln\Lambda\approx 15$)[26].
In many plasma edge scenarios, electrostatic and pressure forces are negligible [27], simplifying equation 4 to focus solely on friction and ion thermal forces, independent of charge state $Z$. At steady-state, this simplifies to a flow velocity balance [27, 17]:
@@ 3 Coupling impurity transport with erosion and deposition dynamics in WallDYN3D
where $m_{ei,wk,wj,qi}$ describes the charge state resolved redistribution and migration matrix, and $R_{ei,wk,qi}$, and $Y_{ei,em,wk,qm}$, describe the reflection and erosion yields, respectively.
The governing equation for the areal density (atoms/m2) time evolution of wall surface element $\delta_{ei,wk}$ is defined as:
@@
Figure 2: (a, b) Distributions of neutral boron ($B^{0}$) and (c-g) ionized boron states ($B^{+1}$ to $B^{+5}$) at the B powder injection location in the torus. (h) Parallel flow velocity profile (Mach number) with blue and red indicating flows toward the inner and outer divertor targets, respectively.
@@ 4 Full-torus modeling of boron powder injection and transport with EMC3-EIRENE and DIS
A representative L-mode scenario for real-time wall conditioning in the DIII-D tokamak has been implemented using the EMC3-EIRENE code. This implementation utilizes a full-torus (360) grid based on the equilibrium #$\#179894$ at 3500 ms, specifically designed to fully capture the migration of boron powder particles and the resulting distributions of boron sources and fluxes. Previously, this scenario was used to model boron transport from a single atomic source [17]. In the current implementation, boron is introduced through powder injection, and the ablation and transport of this boron dust are simulated using the DIS code. A toroidal cross-section of the plasma grid domain for EMC3-EIRENE is shown in figure 1(a). The impurity powder source is situated at the tip of the upper small angle slot (SAS) divertor, where it is connected via a vertical tube to the powder dropper. The simulation accounts for a 5 cm diameter gap at this dropping location in the divertor geometry.
@@
In the first step, the plasma and neutral transport have been calculated for the reference scenario without impurity injection for the entire torus. An upstream density of $n_{up}=1.5\cdot 10^{19}$ m-3 was set based on experimental measurements. The anomalous particle and heat cross-field transport in the plasma edge was prescribed by $D_{\perp}=0.33$ m2s-1 and $\chi_{\perp}=1.0$ m2s-1 for the purpose of this study based on reference values [17]. The cross-field diffusivity of main plasma and impurity ions is assumed to be equal for simplicity.
Figure 1(a,b) displays the boron source distributions along powder particle trajectories for sizes ranging from 5 to 250 $\mu$m, covering typical experimental particle diameters. The evaporation usually occurs when the powder particles reach the separatrix, where the electron temperature and associated sublimation heat flux $Q_{subl}$ increase, increasing the powder particle temperature to the sublimation point. For the mass loss rate follows: $\frac{dm_{d}}{dt}\propto Q_{subl}$. Larger particles, such as those measuring 150 and 250 $\mu$m, follow a predominantly vertical trajectory due to gravity and may cross the separatrix before fully ablating. Conversely, the trajectories of smaller particles (5 $\mu$m) are primarily influenced by the drag and centrifugal forces exerted by the toroidal plasma flow, which pushes these unconfined particles outward.
[…]
A notable increase in boron density is found on the high-field side of the plasma. This inboard peaking of the B density is attributed to the dominant parallel friction force terms associated with the parallel main ion flow, as delineated in equations 2 and 2 and illustrated as Mach flow profile in Figure 2. The parallel main ion flow in the powder deposition region shows a dominant component toward the inboard target (blue region), i.e., pushing the boron impurities effectively poloidally counter-clockwise. Thus, the inward directional bias of the flow culminates in heightened boron densities and fluxes impacting the plasma-facing components on the high-field side.
The boron incident flux distributions on the inner and outer divertor targets reveal a substantial asymmetry, with over $95\%$ of the total boron flux directed towards the inboard target according to the trace impurity fluid model. The toroidally averaged boron fluxes are shown for all particle sizes in Figure 1(c). The particle size does not significantly affect this inboard/outboard flux imbalance, with the frictional forces ($\propto V_{fr}\propto M_{||}$) of the main ion flow primarily governing the transport of boron in the plasma scenario considered.
Figure 4: Comparative analysis of boron migration: (a) fluid model, (b) kinetic model, and (c) numerical differences between fluid and kinetic models.
However, this boron distribution only represents direct deposition from the injection source without considering any subsequent removal or migration upon contact with the target surfaces, i.e., impurities do not recycle. As a result, deposition is narrowly confined to the far SOL, specifically between $Z=-115$ and $-95\,\text{cm}$ on the inboard divertor. Additionally, toroidal asymmetries in the boron deposition on the outer divertor arise from the localized nature of the boron source and the specific orientation of the flux tubes connecting the source region to the divertor targets, as noted in [17].
@@
The source rates were scaled to $10^{20}$ B/s and $10^{21}$ B/s to explore how experimental injection levels of a few milligrams per second up to 20 mg/s may affect the surface distribution and chemical composition. Carbon, the predominant intrinsic impurity in the DIII-D tokamak due to its graphite PFCs, was included in the simulation. The sputtering yields for C, both chemical and charge exchange, were set at 1%, with the physical sputtering yield determined by the energies imparted by the equation $3qT_{e}+2T_{i}$. The chemical and charge exchange sputtering for B were neglected. Multiple WallDYN3D time steps were carried out to model the evolution of the areal densities for B and C on the PFCs within the plasma domain. In each step, equations (5) and (6) are solved, resulting in effective recycling of the impurities.
@@
A 3D visualization of the boron fluxes impacting the divertor targets at an IPD injection rate of $10^{20}$ B/s, as simulated using WallDYN3D, is shown in Figure 6. This figure initially depicts the boron deposition without reflections, erosion, or redeposition (Figure 6(a)). It then progresses through time steps from 0 to 400 seconds, incorporating these effects (Figure 6(b-f)). Notably, after reaching equilibrium after approximately 20 seconds, there is a significant alteration in the boron flux distribution. The outboard divertor target experiences an increased flux compared to the inboard target, as detailed in figures 6(e,f).
@@
In WallDYN3D, impurity transport is treated kinetically by default [28]. A comparison between kinetic and fluid modeling was conducted to investigate potential differences in the transport between the high and low-field sides of relatively light B particles. In this case, boron is sourced from the high field side target. In figure 4, the B flux mappings are shown for the kinetic (Figure 4a) and fluid (Figure 4b) approaches. The boron fluxes are shown as charge state integrated re-distribution matrices, simplifying the three-dimensional migration data into a more comprehensible two-dimensional format through weighted averaging. The diagonal terms indicate the dominating local redeposition of B on the inner and outer divertor targets. The matrices also reveal an absence of transport from the inner to the outer target. However, values in the bottom right sections indicate additional transport from the outer to the inner target. This behavior is consistent in both the fluid and kinetic treatment qualitatively and quantitatively, as the difference in figure 6c shows. Therefore, the remaining differences in the initial deposition between the two models can be mainly attributed to differences in the impurity transport of boron ionized near the injection location.
Figure 6: (a) Boron surface concentrations after 400 iterations for moderate ($10^{20}$ B/s) and high ($10^{21}$ B/s) boron source strengths. (b) XPS measurement depth profile on DiMES sample from similar experiment showing a 40% boron concentration (40% carbon, not shown: 20% oxygen and other impurities) [2].

### v2

@@ Abstract
An integrated modeling framework for investigating the application of solid boron (B) powder injection for real-time surface conditioning of plasma-facing components (PFCs) in tokamak environments is presented. Utilizing the DIII-D impurity powder dropper (IPD) setup, this study simulates B powder injection scenarios ranging from milligrams to tens of milligrams per second, corresponding to boron flux rates of $10^{20}-10^{21}$ B/s in standard L-mode conditions. The comprehensive modeling approach combines EMC3-EIRENE for simulating the deuterium plasma background and the Dust Injection Simulator (DIS) for the ablation and transport of the boron powder particles. The results show substantial transport of boron to the inboard lower divertor, predominantly influenced by the main ion plasma flow. The dependency on powder particle size (5-250 $\mu$m) was found to be insignificant for the scenario considered. The effects of erosion and redeposition were considered to reconcile the discrepancies with experimental observations, which saw substantial deposition on the outer divertor plasma-facing components. For this purpose, the WallDYN3D code was updated to include boron sources within the plasma domain and integrated into the modeling framework. The mixed-material migration modeling shows evolving boron deposition patterns, suggesting the formation of mixed B-C layers or predominantly B coverage depending on the powder mass flow rate. While the modeling outcomes at lower B injection rates tend to align with DIII-D experimental observations, the prediction of near-pure boron layers at higher rates has yet to be experimentally verified in the carbon environment of the DIII-D tokamak. The extensive reach of boron layers found in the modeling suggests the need for modeling that encompasses the entire wall geometry for more accurate experimental correlations. This integrated approach sets a precedent for analyzing and applying real-time in-situ boron coating techniques in advanced tokamak scenarios, potentially extendable to ITER.
@@ 1 Introduction
Real-time wall conditioning (RTWC) has recently gained attention in tokamak and stellarator studies [1, 2, 3, 4, 5, 6, 7, 8, 9]. Traditional techniques, like glow discharge boronization, use hazardous borane gases and interrupt plasma operation, making them unsuitable for long-pulse reactors. They also condition divertor PFCs less effectively than the main chamber walls. For ITER, the strategic decision to replace the beryllium wall with a full tungsten (W) one offers benefits such as better tritium compatibility, lower dust production, and enhanced resilience to thermal loads during disruptions. While current analysis confirms that Q=10 operation is achievable, investigating additional risk-mitigation strategies, such as boron material injection, remains beneficial [10]. Similarly, in the context of the W7-X and DIII-D wall change-out projects [11, 12], which are considering tungsten as a reactor-relevant material, real-time boronization may facilitate high-Z impurity control and support access to low-collisionality plasmas. The development of impurity powder dropper (IPD) technology [13], which releases various impurity powders into the plasma in real time, has significantly advanced RTWC research. The IPD dispenses non-gaseous impurities like lithium, boron, boron nitride, and silicon. Understanding the transport and deposition of these impurities in the scrape-off layer (SOL) is crucial for ITER and future fusion devices.
This paper presents a new framework for modeling solid boron injection, the key material for real-time wall conditioning in tokamaks, stellarators, and heliotrons. It integrates transport, erosion, and deposition modeling to assess boron coverage on divertor PFCs, crucial for evaluating its suitability in future devices like ITER and FPPs [14, 15, 16]. The goal is to demonstrate proof of concept by coupling three models based on a DIII-D RTWC scenario from a prior study [17].
@@
The EMC3-EIRENE code [18, 19] simulates full torus plasma, impurity transport, and plasma-material interactions. It has revealed non-axisymmetric effects from atomic impurity sources in axis-symmetric setups, such as nitrogen seeding in Alcator C-Mod and boron injection in DIII-D [20, 17]. The Dust Injection Simulator (DIS) [21] models dust transport, similar to DUSTT, which has been applied to intrinsic dust in 2D plasma [22, 23] and injected dust in the 3D heliotron LHD [24].
@@
To decipher the mechanisms that lead to B wall coatings through B powder injection, it is crucial to correlate deposition with transport. For this purpose, the WallDYN3D code was employed, which has been recently extended to 3D configurations to enable studies in both tokamaks with non-axisymmetric wall geometries, RMPs, or other symmetry breaking such as local gas puffs and inherently three-dimensional configurations like stellarators [25, 26, 27]. The forthcoming sections of this paper briefly introduce the three models used (Sec. 2 and 3), discuss the transport effects, and resulting B distribution without impurity recycling (Sec. 4), proceed with the analysis based on mixed-material modeling (Sec. 3 and 6), and finally, present a summary and conclusions (Sec. 7).
@@ 2 Coupled dust and plasma impurity transport modeling with DIS and EMC3-EIRENE
The Dust Injection Simulator (DIS) model was designed to study the dynamics of dust particles within three-dimensional, time-evolving plasmas [21]. It employs a comprehensive approach, integrating equations that account for the motion, charge balance, temperature changes, and mass evolution of the dust particles. These equations allow for the simulation of dust transport and ablation processes, including evaporation.
@@
where $R_{f}$ describes the force acting on the powder particles consisting of the drag force exerted by the plasma ($F_{drag}$), electrostatic force ($F_{E}$), gravitational force ($F_{grav}$), and centrifugal force ($F_{centr.}$). Presently, $F_{E}$ does not contribute to the dust transport since EMC3-EIRENE does not yet solve the electric field self-consistently. Particles reaching the wall collide elastically, reflect back, and continue evaporating in the plasma. The deposition depths in the plasma of the injected materials are determined by material-specific properties such as vapor pressure, heat sublimation, emissivity, secondary electron emission, material density, and boiling and melting points. Note that the present version of the model assumes perfect spherical symmetry of particles, neglecting effects such as the rocket force [28].
@@
where $S_{z}$ is the impurity ionization source for charge state $Z$, $D_{imp}$ the anomalous diffusivity, $n_{z}$ and $p_{z}$ the impurity density and pressure, $E_{\parallel}$ the parallel electric field, and $V_{\parallel}$ and $V_{z\parallel}$ the main and impurity ion parallel velocities. The impurity-main ion collision time is given by $\tau_{zi}=\frac{1.47\cdot 10^{13}m_{z}(\nicefrac{{T_{i}^{3}}}{{m_{i}}})^{0.5}}{(1+\nicefrac{{m_{i}}}{{m_{z}}})n_{i}Z^{2}ln\Lambda}$ with $\ln\Lambda\approx 15$ [29]. Equation 3 terms correspond to pressure ($F_{P}$), electric ($F_{E}$), friction ($F_{fr}$), and electron/ion thermal forces ($F_{th,e}$, $F_{th,i}$). The impurity temperature is assumed equal to the main ion temperature ($T_{z}=T_{i}$).
In many plasma edge scenarios, electrostatic and pressure forces can be neglected, reducing equation 3 to focus on friction and ion thermal forces, independent of charge state $Z$. At steady state, this leads to a flow velocity balance [30, 17]:
@@ 3 Coupling impurity transport with erosion and deposition dynamics in WallDYN3D
where $m_{ei,wk,wj,qi}$ describes the charge state resolved redistribution and migration matrix, and $R_{ei,wk,qi}$, and $Y_{ei,em,wk,qm}$, describe the reflection and erosion yields, respectively. The governing equation for the areal density (atoms/m2) time evolution of wall surface element $\delta_{ei,wk}$ is defined as:
@@
Figure 1: (a) Top-down view and (b) cross-sectional view on boron source distribution resulting from ablation of 5-250 $\mu$m size powder particles along the trajectories calculated with DIS. (c) Boron flux profile at the lower inner divertor (peak location) for 5-250 $\mu$m B powder particles (zero impurity recycling), shown along the Z-coordinate of the inner divertor target.
@@ 4 Full-torus modeling of boron powder injection and transport with EMC3-EIRENE and DIS
A representative L-mode deuterium (D) plasma scenario for real-time wall conditioning in the DIII-D tokamak has been implemented using the EMC3-EIRENE code. This implementation utilizes a full-torus (360) grid based on the equilibrium #$\#179894$ at 3500 ms, specifically designed to fully capture the migration of boron powder particles and the resulting distributions of boron sources and fluxes. Previously, this scenario was used to model boron transport from a single atomic source [17]. In the current implementation, boron is introduced through powder injection, and the ablation and transport of this boron dust are simulated using the DIS code. A toroidal cross-section of the plasma grid domain for EMC3-EIRENE is shown in figure 1(a). The impurity powder source is situated at the tip of the upper small angle slot (SAS) divertor, where it is connected via a vertical tube to the powder dropper. The simulation accounts for a 5 cm diameter gap at this dropping location in the divertor geometry.
In the first step, the deuterium plasma and neutral transport have been calculated for the reference scenario without impurity injection for the entire torus. An upstream density of $n_{up}=1.5\cdot 10^{19}$ m-3 was set based on experimental measurements, resulting in recycling flux $\Gamma_{rec}=13.3$ kA. The anomalous particle and heat cross-field transport in the plasma edge was prescribed by $D_{\perp}=0.33$ m2s-1 and $\chi_{\perp}=1.0$ m2s-1 for the purpose of this study based on reference values [17]. The cross-field diffusivity of main plasma and impurity ions is assumed to be equal for simplicity.
Figure 2: (a, b) Distributions of neutral boron ($B^{0}$) and (c-g) ionized boron states ($B^{+1}$ to $B^{+5}$) at the powder injection location in the torus, considering 45 $\mu$m boron particles. (h) Parallel flow velocity profile (Mach number) with blue and red indicating flows toward the inner and outer divertor targets, respectively.
Figure 1(a,b) displays boron source distributions along particle trajectories for sizes from 5 to 250 $\mu$m, representing typical experimental diameters. The injected mass flow rate is constant for all particle sizes. The evaporation usually occurs when the powder particles reach the separatrix, where the electron temperature and associated heat flux $Q_{subl}$ increase, raising the powder particle temperature to the evaporation point. For the mass loss rate follows: $\frac{dm_{d}}{dt}\propto Q_{subl}$. Larger particles, such as those measuring 150 and 250 $\mu$m, follow a predominantly vertical trajectory due to gravity and may cross the separatrix before fully ablating. Conversely, the trajectories of smaller particles (5 $\mu$m) are primarily influenced by the drag and centrifugal forces exerted by the toroidal plasma flow, which pushes these unconfined particles outward. All particle sizes considered in this study evaporate within the simulation domain, with boron ions crossing the inner boundary handled by the 1D core impurity model [18].
[…]
A notable increase in boron density is found on the high-field side of the plasma. This inboard peaking of the B density is attributed to the dominant parallel friction force terms associated with the parallel main ion flow, as delineated in equations 2 and 3 and illustrated as Mach flow profile in Figure 2. The parallel main ion flow in the powder deposition region shows a dominant component toward the inboard target (blue region), i.e., pushing the boron impurities effectively poloidally counter-clockwise. Thus, the inward directional bias of the flow culminates in heightened boron densities and fluxes impacting the plasma-facing components on the high-field side.
@@
The boron incident flux distributions on the inner and outer divertor targets reveal a substantial asymmetry, with over $95\%$ of the total boron flux directed towards the inboard target according to the trace impurity fluid model. The toroidally averaged boron fluxes are shown for all particle sizes in Figure 1(c). The particle size does not significantly affect this inboard/outboard flux imbalance, with the frictional forces ($\propto V_{fr}\propto M_{||}$) of the main ion flow primarily governing the transport of boron in the plasma scenario considered. However, this boron distribution only represents direct deposition from the injection source without considering any subsequent removal or migration upon contact with the target surfaces, i.e., impurities do not recycle. As a result, deposition is narrowly confined to the far SOL, specifically between $Z=-115$ and $-95\,\text{cm}$ on the inboard divertor. Additionally, toroidal asymmetries in the boron deposition on the outer divertor arise from the localized nature of the boron source and the specific orientation of the flux tubes connecting the source region to the divertor targets, as noted in [17]. A limitation of this scenario is that the plasma simulation domain covers only the divertor region, as the plasma grid was originally designed for transport studies rather than plasma-material interaction.
@@
Figure 4: (a) Visualization of the lower divertor region with wall indices (IDX) representing poloidal wall elements. Blue elements are within the simulation domain. Charge state-integrated 2D boron redistribution matrices, as toroidal averages of the 3D matrix, are shown for (b) the fluid model and (c) the kinetic model. ”Source wall IDX” indicates where impurities are eroded, while ”Destination wall IDX” shows where they are deposited.
@@
The source rates were scaled to $10^{20}$ B/s and $10^{21}$ B/s to explore how experimental injection levels of a few milligrams per second up to 20 mg/s may affect the surface distribution and chemical composition (the plasma background was kept constant). Carbon, the predominant intrinsic impurity in the DIII-D tokamak due to its graphite PFCs, was included in the simulation. In the WallDYN3D simulations, the sputtering yields for C, both chemical and charge exchange, were set at 1%. The C flux from CX neutrals was estimated using redistribution matrices informed by EMC3-EIRENE background plasma parameters. Physical sputtering by the background D plasma was also considered, with yields calculated using the energy $3qT_{e}+2T_{i}$ based on the local plasma temperatures. The chemical and charge exchange sputtering for B were neglected. Multiple WallDYN3D time steps were carried out to model the evolution of the areal densities for B and C on the PFCs within the plasma domain. In each step, equations (5) and (6) are solved, resulting in effective recycling of the impurities.
@@
A 3D visualization of the boron fluxes impacting the divertor targets at an IPD injection rate of $10^{20}$ B/s, as simulated using WallDYN3D, is shown in Figure 3. This figure initially depicts the boron deposition without reflections, erosion, or redeposition (Figure 3(a)). It then progresses through time steps from 0 to 400 seconds, incorporating these effects (Figure 3(b-f)). Notably, after reaching quasi-equilibrium after approximately 20 seconds, there is a significant alteration in the boron flux distribution. The outboard divertor target experiences an increased flux compared to the inboard target, as detailed in figures 3(e,f).
@@
In WallDYN3D, impurity transport is treated kinetically by default [31]. A comparison between kinetic and fluid modeling was conducted to investigate potential differences in the transport between the high and low-field sides of relatively light B particles. In this case, boron is sourced from the high field side target. The three-dimensional charge-state integrated boron redistribution matrices have been toroidally averaged, resulting in two-dimensional representations, since this scenario assumed the wall geometry was toroidally symmetric. Figure 4(a) shows the wall elements and corresponding indices (IDX) in the poloidal plane. The boron redistribution matrices are presented for both the kinetic (Figure 4b) and fluid (Figure 4c) approaches. The matrices indicate impurity erosion and deposition locations, specifically on the divertor wall elements. The diagonal terms indicate the dominating local redeposition of B on the inner ($IDX=6-12$) and outer divertor targets ($IDX=19-28$). The matrices also reveal an absence of transport from the inner to the outer target. However, values in the bottom right sections indicate additional transport from the outer to the inner target. This behavior is consistent in both the fluid and kinetic treatment qualitatively and quantitatively. The remaining differences between the kinetic and fluid models stem from how they handle impurity transport near the boron injection location. The fluid model assumes immediate equilibrium with the plasma, simplifying the process and leading to overestimated long-range transport. In contrast, the kinetic model accounts for the time needed for boron ions to accelerate and thermalize, resulting in more accurate, localized deposition.